OpenStudy (anonymous):

A segment has endpoints A (-1, 1) and B (8, 4) . Find the coordinates of the three points that divide the segment into four equal parts. (Named from A to B, X Y Z) (Use fraction format.) X = (, ) Y = (, ) Z = (, )

OpenStudy (shayaan_mustafa):

So easy. look at figure.  For X use mid point formula and use coordinates of A and B, for Y use mid point formula and use coordinates of A and X for Z use mid point formula and use coordinates of X and B midpoint formula is, x=(x1 + x2)/2 and y=(y1 + y2)/2